OpenStudy (anonymous):

1. The volume is found by the function f(x) = x3 + 3x2 – 10x – 24. You need to identify the correct factors of f(x). Suggestions for the factors are: (x – 1) (x + 2) (x – 3) (x + 4) (x + 6) List the correct factors. Then justify your selections with complete sentences. 2. Professor Plum says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. Miss Scarlett she says that 2 is a zero of g(x) because g(2) = 0. Which guest is correct and which guest is incorrect? Explain your answer in complete sentences.

OpenStudy (freckles):

so have you plugged in 1,-2,3,-4,-6 into f to see which would give you 0 as an output

OpenStudy (freckles):

for example f(1)=1^3+3(1)^2-10(1)-24 =1 +3 -10 -24 = 4 - 34=-30 which is not 0 so we know x=1 is not a zero so x-1 is not a factor

OpenStudy (freckles):

what happens when you plug in -2?

OpenStudy (freckles):

if f(n)=0, then x=n is a zero and x-n is a factor

OpenStudy (freckles):

(-2)^3=-8 not -2 (-2)^2=4 not 1 -10(-2)=20 not -10

OpenStudy (freckles):

\[f(-2)=(-2)^3+3(-2)^2-10(-2)-24 \\ \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ }=-8+3(4)+20-24 \\ \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } =-8+12-4=4-4=0 \]

OpenStudy (freckles):

since f(-2)=0 then x=-2 is a zero and x+2 is a factor

OpenStudy (anonymous):

f(3)=(3)^3+3(3)^2-10(3)-24 =27+3(9)-30-24 =27+27-20-24=10

OpenStudy (freckles):

you know you changed -30 to -20

OpenStudy (freckles):

f(3)=54-54=0

OpenStudy (freckles):

so that means x=3 is a zero and x-3 is therefore a factor

OpenStudy (freckles):

so you have f(x)=(x+2)(x-3)*something else

OpenStudy (freckles):

you need to find one more zero

OpenStudy (freckles):

and therefore one more factor

OpenStudy (anonymous):

-4 or -6

OpenStudy (freckles):

plug into f to see which gives you 0

OpenStudy (anonymous):

f(-4)=-4^3+3(-4)^2-10(-4)-24 =-64+48-40-24=-80 so no, x=-4 cant be a factor that leaves -6, and as the last option....it has to be a factor

OpenStudy (freckles):

you are off by a sign in the expression -64+48-40-24

OpenStudy (freckles):

-10(-4)=40

OpenStudy (freckles):

-64+48+40-24=?

OpenStudy (anonymous):

oh, wow....you are a life saver. thanks

OpenStudy (anonymous):

it equals 0

OpenStudy (freckles):

right on! hints for the question: f(n)=0 means x=n is a zero and x-n is a factor f(-n)=0 means x=-n is a zero and x+n is a factor - if we divide something function f(x) by x-r then we have f(x)/(x-r)=q(x)+R/(x-r) where R is the remainder and q(x) is the quotient we can take this same equation and multiply both sides by (x-r) f(x)=q(x)(x-r)+R if we plug r into f we get f(r)=q(r)*(r-r)+R=q(r)*(0)+R=0+R=R so R is the remainder when dividing f(x) by (x-r). Now if R is a zero then x-r is a factor therfore x=r is a zero --- we could have also chose to divide f by x+r so if we did that then we have f(x)=q(x)(x+r)+R pluggining in -r gives f(-r)=R if R again is 0, then x=-r is a zero then x+r is a factor

OpenStudy (freckles):

hints for the next question*